OpenStudy (anonymous):

Suppose $5000 is deposited in a savings account that pays 12% annual interest compounded monthly. (a) What is the monthly interest rate on the account? (b) Assuming that no withdrawals are made. how much money will be in the account after five years? (c) What is the annual yield on this account?

OpenStudy (anonymous):

monthly interest rate = 1% \[A = P(1+r/n)^{nt}\] \[A=5000(1+0.12/12)^{12\times 5}\] \[A=9083.48\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

interest = 4083.48 Annual yield =\[APY = 100 \left[ \left( 1 + 4083.48/5000 \right) ^{1/5} - 1\right]\] APY = 12.68%

OpenStudy (anonymous):

thank you, thank you, thank you!

OpenStudy (anonymous):

medal would help though ;)

OpenStudy (anonymous):

sorry, ive never been on this before :) i think i did it just now.